OpenStudy (anonymous):

Max went to the store and bought three types of apples: Red Delicious for $0.92, Granny Smith for $0.89, and Rome for $1.10. He bought twice as many Red Delicious apples as he did Granny Smith, and he bought one-third as many Rome as he did Red Delicious. How many of each type of apple did he buy if he bought a total of 22 apples? Max went to the store and bought three types of apples: Red Delicious for $0.92, Granny Smith for $0.89, and Rome for $1.10. He bought twice as many Red Delicious apples as he did Granny Smith, and he bought one-third as many Rome as he did Red Delicious. How many of each type of apple did he buy if he bought a total of 22 apples? @Mathematics

OpenStudy (anonymous):

18 Red Delicious: 8 Granny Smith; 6 Rome 12 Red Delicious; 6 Granny Smith; 4 Rome 10 Red Delicious; 5 Granny Smith; 7 Rome 14 Red Delicious; 7 Granny Smith; 3 Rome

OpenStudy (anonymous):

Three types of apples: RD GS R

OpenStudy (anonymous):

He bough twice as many RD as GS. So, RD = 2*GS He bought one-third as many R as he did RD. So, R = 1/3 * RD He bought a total of 22 Apples. So, R + RD + GS = 22. Substitute the first two pieces of information in terms of RD into the 3rd. You get: 1/3*RD + 1/2*RD + RD = 22 => 11/6 * RD = 22 => RD = 22*6/11 = 12 => GS = 6 => R = 4